An electron that has velocity v=(3x10^5 m/s) and moves along positive x direction through the uniform magnetic field B=(0.8T) which is along the positive z direction.

(a) Find the force on the electron (magnitude and direction). (e=1.6x10^-31 kg)

(b) Calculate the radius of the electron's path in the magnetic field.
Answer #1

a. F = qvB

F = 1.6*10^-19 * 3*10^5 * 0.8 = 3.84 *10^-14 N


b. r= mv/qB

r = 9.11*10^-31* 3*10^5/1.6*10^-19 * 0.8

r = 2.13 um
Answer #2

F= qvB

F=1.6 * 10 ^ -19 *3 * 10 ^ 5*0.8=3.84 * 10 ^-14 N

F= m*v ^2 /r

r= m*v ^2/F

here mass of the electron is 9.1 * 10 ^31 Kg

r=81.9 *10 ^10/(3.84 * 10 ^-14)=21.32 *10^4 m